OpenStudy (anonymous):

What's the derivative of (sin(x))^(57)?

OpenStudy (turingtest):

\[\large u=v^{57}\]\[\large v=\sin x\]chain rule is\[\large\frac{du}{dx}=\frac{du}{dv}\cdot\frac{dv}{dx}=57v^{56}\cos x=57\sin^{56}x\cos x\]